OpenStudy (quinnlover):

The sum of three numbers is 26. The second number is twice the first and the third number is 6 more than the second. Find the numbers.

OpenStudy (daniel2raw):

the numbers will be x,2x,2x+6

OpenStudy (quinnlover):

thanks

OpenStudy (daniel2raw):

welcome

OpenStudy (shansha1):

the numbers will be x,2x,2x+6 x+2x+2x+6=26 5x+6=26-6 5x=20 /4 x=4 4,8,14

OpenStudy (quinnlover):

wow thanks!!!

OpenStudy (daniel2raw):

lol i just said that but ok

OpenStudy (quinnlover):

i know and thank you too!